Differential Analysis for a Discontinued Product

A condensed income statement by product line for Warrick Beverage Inc. indicated the following for Mango Cola for the past year:

Sales $236,800
Cost of goods sold (109,000)
Gross profit $127,800
Operating expenses (144,000)
Operating loss $(16,200)

It is estimated that 13% of the cost of goods sold represents fixed factory overhead costs and that 21% of the operating expenses are fixed. Because Mango Cola is only one of many products, the fixed costs will not be materially affected if the product is discontinued.

a. Prepare a differential analysis dated February 29 to determine whether Mango Cola should be continued (Alternative 1) or discontinued (Alternative 2). If an amount is zero, enter "0". If required, use a minus sign to indicate a loss.

Differential Analysis
Continue (Alt. 1) or Discontinue (Alt. 2) Mango Cola
February 29
Continue
Mango Cola
(Alternative 1)
Discontinue
Mango Cola
(Alternative 2)
Differential
Effects
(Alternative 2)
Revenues $ $ $
Costs:
Variable cost of goods sold
Variable operating expenses
Fixed costs
Profit (Loss) $ $ $
Answer #1

Answer-a)-

WARRICK BEVERAGE INC.
DIFFERENTIAL ANALYSIS
PARTICULARS CONTINUE MANGO COLA DISCONTINUE MANGO COLA NET INCOME INCREASE (DECREASE)
$ $ $
Sales 236800 0 -236800
Less- Variable Costs
Cost of goods sold $109000*87%=$94830 0 94830
Operating expenses $144000*79%= $113760 0 113760
Total Variable costs 208590 0 208590
Contribution margin 28210 0 -28210
Less- Fixed costs
Cost of goods sold $109000*13%=$14170 14170 0
Operating expenses $144000*21%= $30240 30240 0
Total Fixed costs 44410 44410 0
Net Income (Loss) -16200 -44410 -28210

Mamgo cola should not be discontinued.